please help me solve this problem
1.a rectangular plot of ground is 120 yards long and 80 yards wide.To double the area of the plot, while rataining the rectangular shape, strips of equal width will be added at one end and along one sde. Find the width of the strip?

Answer by ankor@dixie-net.com(22740)

1.a rectangular plot of ground is 120 yards long and 80 yards wide.To double the area of the plot, while retaining the rectangular shape, strips of equal width will be added at one end and along one side. Find the width of the strip?
:
Let x = width of the strip
:
Dimensions of new rectangle will be (x+120) by (x+80)
Old rectangle area = 120*80 = 9600
:
Write an area equation for the new rectangle
(x+120)*(x+80) = 2(9600)
FOIL
x^2 + 80x + 120x + 9600 = 19200
:
x^2 + 200x + 9600 - 19200 = 0
:
x^2 + 200x - 9600 = 0
Factors to:
(x+240)(x-40) = 0
Positive solution
x = +40 yds is the width of the strip that doubles the area
;
;
Check solution:
(40+120) * (40+80) = 19200 which is double the original area